n2-1 Find the sum of the series (n=1 n=1 (n2 +1)2
Answer #1

x n2 – 1 calculate Σ n=1 (n2 + 1)? Simplify: 8W n2 – 1 = ♡ n=1 (n2 + 1)2 (n − 1) (n + 1) (n2 + 1)? n=1 Perform partial fractiThe form of the partial fraction decomposition is n2 1 An + B n2 + 1 + Cn + D (n2 + 1) (n2 + 1)? Write the right-hand side asCollect up the like terms: na – 1= n’ A+n+B+n (A + C) + B + D The coefficients near the like terms should be equal, so the fo(n − 1) (n +1) (n2 + 1)? m*) - (( +1 4w;)) n=1 Split the series: 1 2 (n2 +1) 2 (n2 + 1)2 1 n=1 Split the series: CO 0 1 1 1Pull the constant out of the series: 1 + a coth () 2 + 2 n=1 2(n − 1)2 + Σ + n=o 2(n +1+i)? i coth (T) + 2n - 2i 2 2 i i 1 (nTherefore, 00 1 n=0 i i + (n+1-i) 2 + a coth (T) 2 + 2 2n + 2i 2n - 2i n=1 Ś (2,1+i) + 2 Ś (2,1+i) + 2 0 1 i i (5(2,1 - i)) 2